Computer screen (laptop)
Need to find the length of diagonal
Height = 9.5"
Length = 14"
Width = .5"
What is the length of diagonal?
When I use a tape measure, it comes to 17", but it's not working out in the solution. Why?

Answer by rfer(16322)

Your measuring tape is close.
The answer is 16.9
a^2+b^2=c^2
9.5^2+14^2=c^2
90.25+196=286.25
sq rt of 286.25=16.9
----------------------------
the formula for solving for the hypotenuse is the Pythagorean theorem
which is side "a" squared plus side "b" squared equals the hypotenuse squared, which is side "c". So then you need to take the sq root of the hypotenuse squared for the answer.
---------------------------------------
a^2 is the way we write "a" squared on the computer. That is where the 2's came from.
Does that help?
Good luck
Bob